### v1

# NoisyGRPO: Incentivizing Multimodal CoT Reasoning via Noise Injection and Bayesian Estimation

Abstract
Reinforcement learning (RL) has shown promise in enhancing the general Chain-of-Thought (CoT) reasoning capabilities of multimodal large language models (MLLMs). However, when applied to improve general CoT reasoning, existing RL frameworks often struggle to generalize beyond the training distribution. To address this, we propose NoisyGRPO, a systematic multimodal RL framework that introduces controllable noise into visual inputs for enhanced exploration and explicitly models the advantage estimation process via a Bayesian framework. Specifically, NoisyGRPO improves RL training by: (1) Noise-Injected Exploration Policy: Perturbing visual inputs with Gaussian noise to encourage exploration across a wider range of visual scenarios; and (2) Bayesian Advantage Estimation: Formulating advantage estimation as a principled Bayesian inference problem, where the injected noise level serves as a prior and the observed trajectory reward as the likelihood. This Bayesian modeling fuses both sources of information to compute a robust posterior estimate of trajectory advantage, effectively guiding MLLMs to prefer visually grounded trajectories over noisy ones. Experiments on standard CoT quality, general capability, and hallucination benchmarks demonstrate that NoisyGRPO substantially improves generalization and robustness, especially in RL settings with small-scale MLLMs such as Qwen2.5-VL 3B. The project page is available at https://artanic30.github.io/project_pages/NoisyGRPO.
1 Introduction
The success of OpenAI’s O1 [openai2024o1] and DeepSeek-R1 [DeepSeekR1] highlights that activating Chain-of-Thought (CoT) reasoning abilities through Reinforcement Learning (RL) is a promising strategy for scaling model intelligence at test time. In particular, value-free RL methods such as Group Relative Policy Optimization [GRPO] (GRPO) have demonstrated substantial improvements in LLM performance while avoiding the costly requirement of training a separate value model. As recent works show [Peng2025LMMR1E3, Huang2025VisionR1IR, Deng2025OpenVLThinkerAE, Meng2025MMEurekaET, vlmr1], this efficiency in both data and computation makes GRPO an appealing approach for extension to complex visual language tasks such as visual math problem solving.
However, when applying GRPO to improve general chain-of-thought (CoT) reasoning in multimodal large language models (MLLMs), we observe notable limitations in the model’s ability to generalize beyond the training distribution, which hinders the broad applicability of MLLMs. As shown in Figure 1, higher training rewards under the GRPO framework do not consistently yield better evaluation performance. We identify two primary factors contributing to this issue: (1) Limited Policy Exploration. RL in the generative model domain utilizes temperature sampling to generate multiple rollouts for policy exploration; however, as noted in DAPO [DAPO], these rollouts often converge to near-identical outputs during training, leading to insufficient exploration and early deterministic policy. (2) Missing process supervision. Rule-based rewards judge only the final answer, leaving the intermediate reasoning steps unsupervised; the policy, therefore, learns shortcuts and visual hallucinations that do not transfer out of distribution.
Figure 1: Performance and training statistics of three RL methods. GRPO with noise injection refers to GRPO trained with noise-perturbed rollouts. The left plot shows evaluation performance on the MMStar benchmark over training iterations. The middle plot presents the standard deviation of rewards, reflecting the exploration degree of the policy. The right plot shows the accuracy reward, indicating how well the model fits the training data. Shaded areas represent the variance of the corresponding metrics.
To address these challenges, we propose  NoisyGRPO, a unified framework that enhances multimodal reinforcement learning, such as GRPO [GRPO] by both promoting thorough exploration and ensuring robust policy optimization. Our key innovation is to inject controllable noise into visual inputs during rollouts, thereby encouraging diverse exploration and providing an unbiased, trajectory-level measure of reasoning difficulty. However, while promoting exploration, noise injection induces an off-policy scenario: the behavior policy that collects trajectories generated with noise-disturbed images differs from the target policy that operates on clean visual input. This distribution shift causes the value function to be queried on out-of-distribution states, which introduces bias into the advantage estimates and corrupts the policy gradient, as illustrated in Figure 1. To tackle this issue, we introduce a Bayesian advantage estimation framework that treats the injected noise as a prior and the trajectory reward as the likelihood within a principled Bayesian formulation, which allows us to explicitly calibrate the policy update according to trajectory-level noise and reward signals.
NoisyGRPO consists of two key components without additional computational overhead. First, Noise Injection augments policy exploration by applying Gaussian noise of various magnitudes (sampled from a uniform distribution) to input images in rollouts, maintaining trajectory diversity (indicated by the consistently high Reward Std shown in Figure 1) and reducing hallucinations via implicitly favoring visually grounded reasoning. The second component is Bayesian Advantage Estimation, which integrates the magnitude of the injected noise with observed trajectory rewards to yield more accurate advantage estimates. Specifically, we treat the likelihood uncertainty as a fixed constant while modeling the prior uncertainty as a function of the variance among rewards within the same trajectory group. The intuition is that if trajectory rewards are similar despite varying noise magnitudes, the prior uncertainty remains high; in contrast, larger reward variation across noise levels corresponds to reduced prior uncertainty. By defining advantage estimation in this Bayesian manner, we obtain a posterior that adaptively incorporates both noise magnitude and trajectory reward statistics, leading to more accurate and robust advantage estimates.
To demonstrate NoisyGRPO’s effectiveness, we evaluate its impact on improving MLLMs’ general capabilities, which is a more challenging setting than addressing domain-specific problems. Specifically, we evaluate NoisyGRPO from three dimensions: CoT quality [mmecot], comprehensive MLLM capability [mmstar, mmerealworld], and hallucination [amber]. Results show that NoisyGRPO consistently outperforms GRPO, particularly with small-scale MLLM—for instance, achieving a +4.4 improvement in CoT quality on MME-COT and a +3.7 gain in average performance on MMStar with Qwen2.5-VL 3B.
Our main contributions are summarized as follows:
  • We propose NoisyGRPO, a reinforcement learning framework designed to enhance the multimodal Chain-of-Thought (CoT) reasoning capabilities of MLLMs on general tasks.
  • NoisyGRPO employs a noise injection strategy to promote policy exploration, and introduces a Bayesian advantage estimation method that mitigates the negative effects of noise by combining prior estimates from noise magnitude with observed response correctness.
  • Experimental results show that NoisyGRPO consistently outperforms GRPO across CoT quality, general capability, and hallucination benchmarks, with especially strong gains in RL with small-scale MLLMs.
2 Related Works
Multi-modal Large Language Model
Vision-language models (VLMs) have rapidly advanced in their ability to comprehend and reason over both visual and textual modalities and demonstrate great improvements over downstream tasks [oquab2023dinov2, cocoop, guo2022calip, Feng2022PromptDetTO, HOICLIP, maccap]. These models typically integrate visual encoders with large language models (LLMs) to enable cross-modal understanding and inference. Foundational works such as Flamingo [alayrac2022flamingo] and BLIP-2 [li2023blip] established effective strategies for aligning vision and language components. The LLaVA-series [Liu2023VisualIT, liu2024llavanext, Li2024LLaVAOneVisionEV] and SPHINX-series [SPHINX, SPHINX-X] further advanced the field by introducing visual instruction tuning, significantly enhancing multimodal capabilities. Large-scale models like GPT-4o [Hurst2024GPT4oSC] and Gemini [Reid2024Gemini1U] have demonstrated strong general-purpose visual understanding through large-scale multimodal pertaining. To address scalability, mixture-of-experts approaches—such as DeepSeek-VL2 [Wu2024DeepSeekVL2MV], Uni-MoE [Li2024UniMoESU], and MoVA [Zong2024MoVAAM]—improve computational efficiency by selectively activating expert modules based on input characteristics. More recently, domain-specific methods such as Math-LLaVA [Shi2024MathLLaVABM] and MAVIS [Zhang2024MAVISMV] employed mathematical visual instruction tuning to improve VLMs’ ability to interpret and solve complex multimodal math problems. At the same time, unified architectures like SEED-X [Ge2024SEEDXMM], Chameleon [Team2024ChameleonME], Show-O [Xie2024ShowoOS], and the Janus series [Wu2024JanusDV, Ma2024JanusFlowHA, Chen2024ExpandingPB] integrate both visual understanding and generation capabilities within a single framework. Despite these advances, most existing VLMs still struggle with robust visual reasoning, particularly in tasks that demand deep visual analysis and complex multi-step reasoning.
Reinforcement Learning in LLMs and VLMs
Reinforcement Learning (RL) has emerged as a key technique for enhancing the capabilities of Large Language Models (LLMs) and Multimodal Large Language Models (MLLMs). Early advancements focused on Reinforcement Learning from Human Feedback (RLHF) [Ouyang2022TrainingLM], which aligns model outputs with human preferences [Achiam2023GPT4TR]. More recent research has demonstrated that RL-based methods can significantly improve reasoning abilities. For instance, DeepSeek-R1 [DeepSeekR1] integrates rule-based rewards with Group Relative Policy Optimization (GRPO) [Shao2024DeepSeekMathPT], while Kimi-1.5 [Team2025KimiKS] employs a variant of online policy mirror descent—both achieving notable gains in reasoning performance. In the multimodal domain, RL has also shown promise. Direct Preference Optimization (DPO) [Rafailov2023DirectPO], a simple yet effective RL method, has been applied to reduce hallucination in MLLMs [BPO, Ouali2024CLIPDPOVM, Zhang2025MMRLHFTN]. GRPO has also been extended to vision-language models for specialized tasks such as geometry reasoning and object counting [Peng2025LMMR1E3, Huang2025VisionR1IR, Deng2025OpenVLThinkerAE, Meng2025MMEurekaET, vlmr1]. Though these works apply the GRPO framework to multimodal tasks, their focuses differ from ours. For example, VLM-R1 [vlmr1] adapts it to detection tasks with task-specific rewards. In contrast, we propose a principled framework that systematically improves multimodal RL algorithms like GRPO in multimodal settings by enhancing both policy exploration and advantage estimation.
Figure 2: Overall pipeline of NoisyGRPO. For each image-question pair, we sample noise and inject it into the image. The policy model generates rollouts based on the perturbed inputs, and the reward function evaluates them. We then compute the posterior advantage by combining the noise-based prior with the reward-based observation.
3 NoisyGRPO
In this section, we introduce NoisyGRPO, a reinforcement learning framework designed to enhance the Chain-of-Thought reasoning capabilities of multimodal large language models (MLLMs). To address the generalization limitations of vanilla GRPO when applied to MLLMs, NoisyGRPO encourages policy exploration via perturbing visual inputs and mitigates the resulting adverse effects through Bayesian advantage estimation. Specifically, in section 3.1, we present the training objective of NoisyGRPO and highlight its improvements over the standard GRPO. In section 3.2, we introduce the exploration policy with injection noise on the visual input strategy. In section 3.3, we describe the Bayesian advantage estimation method, which incorporates both the level of noise and the trajectory reward to yield more accurate policy gradients. We provide an overall pipeline in Figure 2.
3.1 Training Objective and Reward Function
Training Objective
To highlight the key innovations of NoisyGRPO, we start by presenting its training objective and comparing it with vanilla GRPO [GRPO]. NoisyGRPO differs in two main aspects: (1) Exploration Policy. We introduce visual noise during sampling to encourage more diverse trajectory exploration. (2) Advantage Estimation. By modeling advantage estimation as a Bayesian inference problem, we combine the noise level and observed trajectory reward to compute a more accurate advantage for each trajectory.
Specifically, given a sample $(q,a,I)$ drawn from the dataset $\mathcal{D}$—where $q$ is the input query, $a$ is the ground-truth answer, and $I$ is the corresponding image—we sample $G$ rollouts $\{o_{i}^{n}\}_{i=1}^{G}$ using a noise-injected exploration policy $\pi^{\text{noise}}$, where each $o_{i}^{n}$ is sampled under a different noise level applied to the image. The corresponding group-wise advantage $\tilde{A}^{k}_{i}$ is estimated based on both trajectory reward and noise level. The training objective of NoisyGRPO is defined as follows:
$J_{\text{NoisyGRPO}}(\theta)=\mathbb{E}_{(q,a,I)\sim\mathcal{D},\ \{o_{i}^{n}\}_{i=1}^{G}\sim\bm{\pi^{\text{noise}}_{\theta_{\text{old}}}}(\cdot\mid q,I)}$ (1)
$\Bigg[\frac{1}{G}\sum_{i=1}^{G}\min\Big(\frac{\pi_{\theta}(o_{i}^{n}\mid q,I)}{\pi_{\theta_{\text{old}}}(o^{n}_{i}\mid q,I)}\bm{\tilde{A}^{k}_{i}},\ \text{clip}\Big(\frac{\pi_{\theta}(o_{i}^{n}\mid q,I)}{\pi_{\theta_{\text{old}}}(o_{i}^{n}\mid q,I)},1-\epsilon,1+\epsilon\Big)\ \bm{\tilde{A}^{k}_{i}}\Big)-\beta D_{\text{KL}}(\pi_{\theta}\|\pi_{\text{ref}})\Bigg],$
where $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the current and previous policies respectively, $\epsilon$ is the PPO clipping parameter, $\beta$ is the KL penalty coefficient, $\pi_{\text{ref}}$ is a reference policy, and $D_{\text{KL}}$ denotes KL divergence. The modifications from vanilla GRPO are marked in bold—specifically, the exploration policy $\bm{\pi^{\text{noise}}}$ and the group-relative advantage $\bm{\tilde{A}^{k}_{i}}$. We elaborate on these components in the following sections.
Reward Function
For the reward function, we adopt two types of rule-based rewards inspired by GRPO: Accuracy and Format rewards. The Correctness reward measures whether the model’s final answer is accurate, regardless of the correctness of intermediate CoT steps. The Format reward encourages the model to produce answers in the expected CoT style. It is worth noting that our training data contains many open-ended VQA samples to improve general CoT ability in multimodal settings. Unlike code or math tasks with short, matchable answers, these samples require a more flexible reward. We use a text embedding model such as SBERT [SBERT] to assess semantic similarity for correctness. Although this may introduce embedding bias, our advantage estimation 3.3 helps mitigate it. In conclusion, the reward is defined as follows:
  • Accuracy rewards: We utilize the SBERT model to compute embeddings for both the predicted and ground-truth answers, and define their similarity as the accuracy reward. To filter low-confidence cases, we set the reward to 0 when the similarity is below $\tau$. For yes/no and multiple-choice questions, we apply exact match: a correct match receives a reward of 1, otherwise 0.
  • Format rewards: To ensure CoT reasoning, we introduce a format reward that requires the model’s reasoning process to appear between <think> and </think> tags.
3.2 Noise Injected Exploration Policy
To enhance the generalization of the multimodal RL framework, we introduce a noise-injected exploration strategy to construct our exploration policy $\pi^{\text{noise}}$, as shown in Figure 2. Different from injecting noise into the entire policy model as in [Igl2019GeneralizationIR], we adopt a targeted and structured strategy by applying multi-step Gaussian noise to the input image $I$ based on a diffusion-style forward process. For each rollout, we first sample a noise level $n_{i}\sim\mathcal{U}(0,1)$, representing a normalized noise step. We then perturb $I$ using the corresponding diffusion process at timestep $\lfloor n_{i}\cdot T\rfloor$, where $T$ is the total number of diffusion steps. $n_{i}=0$ yields the original clean image, while $n_{i}=1$ corresponds to the image at the final diffusion step, effectively resembling pure Gaussian noise. This produces a noisy image $I_{n_{i}}$ which is then used to generate the rollout $o_{i}^{n}$ under the exploration policy $\pi^{\text{noise}}$, resulting in a diverse set of responses $\{o_{i}^{n}\}_{i=1}^{G}$ conditioned on different noise intensities. Notably, noise injection is applied only during rollout collection to encourage exploration. Clean inputs are used during policy updates to ensure stable learning and preserve the accuracy of the learned policy $\pi_{\theta}$.
3.3 Bayesian Advantage Estimation
Bayesian Modeling of Trajectory Quality
After collecting rollouts using the noise-injected exploration policy, we formulate the trajectory quality as a latent variable $r_{i}$, which represents the underlying quality of a sampled rollout $o_{i}$ obtained under noise-injected exploration. To integrate both prior knowledge from noise level and observation, such as the trajectory reward, we adopt a Bayesian modeling approach, treating $r_{i}$ as a Gaussian random variable. Specifically, we assume:
$r_{i}\sim\mathcal{N}(\mu_{i},\sigma^{2}),$ (2)
where $\mu_{i}$ denotes the mean of the trajectory quality $r_{i}$ and $\sigma^{2}$ quantifies the uncertainty. To update this prior based on observed evidence, we consider a Bayesian inference framework. Let $\tilde{r}_{i}$ be an observed variable that serves as a noisy indicator of the true quality $r_{i}$. We assume the observation model is Gaussian:
$\tilde{r}_{i}\sim\mathcal{N}(r_{i},\sigma_{s}^{2}),$ (3)
where $\sigma_{s}^{2}$ captures the reliability of the observation process. Under this setup, the posterior distribution of $r_{i}$ given $\tilde{r}_{i}$ remains Gaussian, and the posterior mean can be derived analytically as:
$\hat{r}_{i}=\tilde{r}_{i}+\frac{\sigma_{s}^{2}}{\sigma^{2}+\sigma_{s}^{2}}(\mu_{i}-\tilde{r}_{i}).$ (4)
This posterior mean balances the prior distribution and the observation, weighted inversely by their respective variances. Intuitively, more confident observations (small $\sigma_{s}^{2}$) shift the posterior closer to $\tilde{r}_{i}$, while a more confident prior (small $\sigma^{2}$) keeps the posterior closer to $\mu_{i}$.
Advantage Estimation
For a set of samples $\{(n_{i},o_{i},r_{i}^{s})\}_{i=1}^{G}$, where $n_{i}$ is the injected noise level, $o_{i}$ is the generated trajectory, and $r_{i}^{s}$ is the trajectory semantic reward, which combines both accuracy and format reward. For the prior estimation of trajectory quality $r_{i}$, we set the noisy reward as $r_{i}^{n}=1-n_{i}$ and treat it as the prior estimation. The observation is the semantic reward $r_{i}^{s}$. We apply normalization to make the prior and semantic rewards comparable across the group of $G$ samples. Given a set of scalar values $\{x_{j}\}_{j=1}^{G}$, the normalization operation is defined as:
$\mathrm{Norm}(x_{i},\{x_{j}\}_{j=1}^{G})=\frac{x_{i}-\mathrm{mean}(\{x_{j}\}_{j=1}^{G})}{\mathrm{std}(\{x_{j}\}_{j=1}^{G})},$ (5)
We apply this normalization to both the prior and semantic rewards as follows:
$\hat{r}^{n}_{i}=\mathrm{Norm}(r^{n}_{i},\{r^{n}_{j}\}_{j=1}^{G}),\quad\hat{r}^{s}_{i}=\mathrm{Norm}(r^{s}_{i},\{r^{s}_{j}\}_{j=1}^{G}),$ (6)
For clarity, we assume both normalized signals are corrupted by a Gaussian, which serves as a core assumption in our Bayesian estimation framework. An analysis of this assumption and its empirical justification is provided in the Appendix. Accordingly, we model the latent reward $r_{i}$ as a random variable drawn from a Gaussian prior centered at $\hat{r}^{n}_{i}$, while the observed semantic reward $\hat{r}^{s}_{i}$ is treated as a noisy observation of $r_{i}$:
$r_{i}\sim\mathcal{N}(\hat{r}^{n}_{i},\sigma_{n}^{2}),\qquad\hat{r}^{s}_{i}\sim\mathcal{N}(r_{i},\sigma_{s}^{2}).$ (7)
To this end, the key challenge is to determine the uncertainty of both signals, which is represented by the $\sigma_{n}^{2}$ and $\sigma_{s}^{2}$. We fix $\sigma_{s}^{2}$ to a constant $\alpha$, as the semantic reward is computed using a text embedding model whose bias is relatively stable. To modulate the prior uncertainty, we define the prior variance $\sigma_{n}^{2}$ to be inversely related to the variability of semantic rewards within the group. The underlying intuition is that $\hat{r}^{s}_{i}$ and $\hat{r}^{n}_{i}$ are positively correlated, as both reflect the underlying trajectory quality, and $n_{i}$ is sampled from a uniform distribution. Therefore, when the group-wise variance of $\hat{r}^{s}_{i}$ is small, the certainty of the prior distribution is low. Specifically:
$\sigma_{s}^{2}=\alpha,\qquad\sigma_{n}^{2}=\frac{\gamma}{\gamma+(\mathrm{std}(\{r^{s}_{i}\}_{i=1}^{G}))^{2}},$ (8)
where $\gamma$ is a scale hyperparameter that modulates the influence of semantic reward variability on prior uncertainty. This formulation implies that when the semantic reward varies widely (i.e., larger standard deviation), the prior distribution is considered more reliable and thus assigned a smaller variance. Given the prior and observation, we apply a Bayesian update to obtain the posterior mean reward:
$\hat{r}_{i}=\hat{r}^{s}_{i}+\frac{\sigma_{s}^{2}}{\sigma_{n}^{2}+\sigma_{s}^{2}}(\hat{r}^{n}_{i}-\hat{r}^{s}_{i}).$ (9)
Finally, to compute the trajectory-wise relative advantage within each group, we normalize the posterior reward:
$\tilde{A}^{k}_{i}=\mathrm{Norm}(\hat{r}_{i},\{\hat{r}_{j}\}_{j=1}^{G}).$ (10)
4 Experiments
Table 1: Results on CoT quality benchmark MME-CoT. We report results across three evaluation dimensions of CoT quality to comprehensively characterize the performance of NoisyGRPO.
CoT quality CoT Robustness CoT Efficiency
F1 Precision Recall Avg. Stability Efficacy Avg. Rel. Rate Ref. Quality
LLaVA-OV-7B 30.9 50.9 22.2 -3.4 -3.8 -3.0 91.5 83.0 100.0
LLaVA-OV-72B 36.3 57.3 26.6 -0.2 0.3 -0.6 95.4 90.8 100.0
MiniCPM-V-2.6 39.8 57.3 30.5 -3.5 -4.8 -2.2 92.8 85.7 100.0
InternVL2.5-8B-MPO 43.0 60.4 33.4 0.6 0.3 0.9 94.7 89.3 100.0
Qwen2-VL-72B 56.2 77.3 44.2 -2.1 -6.5 2.4 96.5 92.9 100.0
GPT-4o 64.0 85.4 51.2 2.1 -1.0 5.1 96.0 92.0 100.0
Kimi k1.5 64.2 92.0 49.3 1.4 2.9 0.0 82.2 92.2 72.2
[…]
GRPO 36.2 53.0 27.4 -7.0 -7.5 -6.5 51.1 76.3 26.0
[…]
GRPO 46.8 64.2 36.9 1.6 3.8 -0.6 66.4 87.2 45.7
[…]
4.1 Training Data
To enhance the general Chain-of-Thought (CoT) reasoning ability of MLLMs through reinforcement learning, we adopt the visual question answering (VQA) portion of the MM-RLHF [MMRLHF] training set as our training data. This dataset spans diverse domains, including conversations, safety, multiple-choice questions, captions, and commonsense reasoning. MM-RLHF applies clustering and filtering techniques to curate a high-quality visual instruction-following dataset. In total, it contains 13k VQA samples: 1.2k yes/no questions, 1.3k multiple-choice questions, and 10k open-ended VQA questions. Notably, most of the training data lack structured answers; therefore, we use a text embedding model to compute the accuracy reward. While this may introduce some model bias, our Bayesian advantage estimation helps mitigate its impact.
4.2 Implementation Details
We implement the NoisyGRPO based on the VLM-R1 [vlmr1] reinforcement learning training framework. For the policy model, we choose the state-of-the-art MLLMs Qwen2.5-VL [Qwen2.5-VL] 3B and 7B. For hyperparameters, we follow the default settings provided by VLM-R1 except for the following changes. The number of sampled rollouts $G$ is set to 4 due to computational resource considerations. In noise injection, the upper bound $\sigma$ for the $\mathcal{U}(0,\sigma)$ is set to 1. The $\alpha$ and $\gamma$ are set to 0.1 and 0.01. The threshold $\tau$ for the accuracy reward is set to 0.6. As there is no validation set, we choose all the hyperparameters based on the results on MMStar [mmstar]. The training takes 6 hours for the 3B variant and 7 hours for the 7B on a single node with 8A100 GPUs.
4.3 Evaluation Benchmarks
To comprehensively evaluate the CoT reasoning capabilities of NoisyGRPO, we select benchmarks from three perspectives: (1) CoT Quality Evaluation. We use the MME-CoT [mmecot] benchmark, which assesses the quality, robustness, and efficiency of multi-modal CoT reasoning. The benchmark covers six categories of questions—including math, OCR, logic, science, space-time, and general scene—providing a holistic view of CoT performance. (2) General Capability Evaluation. We adopt the MMStar [mmstar] benchmark, which is constructed by carefully selecting high-quality samples from existing datasets such as AI2D [Kembhavi2016ADI], SEED-Bench [Li2023SEEDBenchBM], ScienceQA [scienceqa], MMBench [liu2023mmbench], MathVista [Lu2023MathVistaEM], and MMMU [mmmu]. By removing low-quality samples, MMStar provides a cleaner and more comprehensive testbed across six reasoning domains. (3) Hallucination Evaluation. We employ the AMBER [amber] benchmark to evaluate both the generative and discriminative hallucination behaviors of the model, offering a thorough analysis of output reliability. Additionally, we include MME-RealWorld-Lite [mmerealworld] to assess VQA performance in real-world scenarios.
4.4 Baselines
To fairly demonstrate the effectiveness of NoisyGRPO, we compare it against strong baselines across all evaluation settings. For the CoT quality benchmark, we compare NoisyGRPO with the GRPO [GRPO] algorithm under different scales of Qwen2.5-VL [Qwen2.5-VL], and also report results from a wide range of MLLMs [Li2024LLaVAOneVisionEV, MiniCPMV, Chen2024ExpandingPB, Qwen2-VL, GPT-4o, Team2025KimiKS, llava1.5, Qwen-VL, CogVLM, InternLM-XComposer2, liu2024llavanext, openai2023gpt4v, Qwen2.5-VL], from 7B to 72B models, as well as proprietary models like GPT-4o [GPT-4o] and Kimi K1.5 [Team2025KimiKS], for reference. For the general reasoning and hallucination benchmarks, in addition to comparisons between NoisyGRPO and GRPO, we include SFT (Supervised Fine-Tuning) baselines to assess the quality and impact of training data. We also report the performance of state-of-the-art MLLMs for broader comparison.
Table 2: Performance Comparison on General, Hallucination, and Real-World Benchmarks. The CP, FP, IR, LR, ST, MA, and AVG denote Coarse Perception, Fine-grained Perception, Instance Reasoning, Logical Reasoning, Science & Technology, Math, and the average performance, respectively. The symbols $\uparrow$ and $\downarrow$ indicate that higher or lower values are preferred. AMB. Gen. and AMB. Dis. represent the generative and discriminative components of the AMBER benchmark.
MMStar $\mathrm{AMB.}^{G}$ $\mathrm{AMB.}^{D}$ MMERW
CP. FP. IR. LR. ST. MA. Avg. Cs $\downarrow$ Cov. $\uparrow$ Hal. $\downarrow$ Cog. $\downarrow$ F1 $\uparrow$ Acc
LLaVA-1.5 7B 58.8 24.0 38.8 24.0 13.6 22.8 30.3 7.8 51.0 36.4 4.2 74.7 -
Qwen-VL-Chat 59.6 32.0 50.8 29.2 22.0 31.6 37.5 5.5 49.4 23.6 1.9 - -
CogVLM-Chat 66.8 36.8 49.2 31.2 23.6 11.6 36.5 5.6 57.2 23.6 1.3 - -
InternLM-XComposer2 70.8 48.8 65.2 56.4 42.0 49.2 55.4 - - - - - -
LLaVA-Next 34B 66.4 52.0 62.4 46.0 32.4 53.6 52.1 - - - - - -
GPT4V 76.6 51.4 66.6 55.8 42.6 49.8 57.1 4.6 67.1 30.7 2.6 - -
[…]
Qwen2.5-VL 3B 70.3 49.1 59.6 55.1 38.5 59.7 55.4 7.6 69.9 55.7 5.8 89.6 42.1
+ SFT 70.3 49.6 59.0 55.4 39.2 60.4 55.7 6.9 70.4 48.8 3.9 88.8 44.0
+ GRPO 68.0 43.3 60.4 57.0 40.0 58.0 54.5 6.7 68.5 44.6 4.2 89.2 40.8
[…]
Qwen2.5-VL 7B 73.3 60.1 72.2 62.0 44.6 67.3 63.2 4.8 63.6 27.5 1.6 87.4 43.6
+ SFT 71.8 58.8 70.2 60.8 45.2 67.4 62.4 5.4 64.9 32.4 2.1 87.5 43.5
+ GRPO 73.7 57.2 70.5 64.2 44.1 65.7 62.6 4.9 64.3 28.9 1.7 88.0 42.3
[…]
4.5 Performance Analysis
Benchmark Results
For the CoT quality evaluation, we report the results of NoisyGRPO and baselines on the MME-CoT [mmecot] benchmark in Table 4. We observe that NoisyGRPO consistently outperforms GRPO on most CoT quality metrics, with particularly significant improvements for the 3B model. Notably, NoisyGRPO also enhances CoT robustness—for example, NoisyGRPO-7B achieves higher robustness scores than proprietary models like GPT-4o [GPT-4o] and Kimi K1.5 [Team2025KimiKS]. However, we observe that vanilla GRPO surpasses NoisyGRPO in CoT efficiency when using Qwen2.5-VL 7B. We attribute this to NoisyGRPO’s design, which prioritizes grounded reasoning over reasoning efficiency.
For the general reasoning and hallucination benchmarks, Table 4.4 presents a comparison between NoisyGRPO and baseline methods. The results demonstrate that NoisyGRPO effectively enhances MLLM performance across multiple domains by strengthening CoT reasoning, achieving notable gains such as +2.8 and +2.6 in average MMStar [mmstar] scores over the base model. However, we observe a performance drop in the Coarse Perception category for both model variants. We attribute this to the injected noise being insufficient to meaningfully perturb the visual cues necessary for coarse-grained perception tasks, leading to suboptimal performance for this specific task. Furthermore, we observe that NoisyGRPO performs worse on the Cov. metric of the AMBER [amber] benchmark, suggesting that the generated captions do not fully cover all elements of the image. We attribute this to NoisyGRPO’s emphasis on grounded responses, which tends to produce more concise answers and, consequently, reduces overall coverage.
We also find that supervised fine-tuning (SFT) on the current 13k dataset fails to bring consistent performance gains, highlighting that the improvements stem from reinforcement learning rather than high-quality data alone. Furthermore, the performance improvements on MME-RealWorld-Lite [mmerealworld] further demonstrate the potential of NoisyGRPO for real-world applications.
Performance Over Training Iterations
In addition to reporting the best model performance across various benchmarks, we further analyze the training dynamics of NoisyGRPO to provide a more comprehensive understanding of its behavior. Specifically, we examine the relationship between training iterations and performance on the MMStar benchmark, which evaluates MLLMs across six distinct reasoning dimensions. As shown in Figure 3, we compare NoisyGRPO with the baseline GRPO in terms of performance evolution, as well as the progression of the importance weight—defined as $\frac{\sigma_{s}^{2}}{\sigma_{n}^{2}+\sigma_{s}^{2}}$ in Eq. 9—throughout training. We also report the average completion length produced by each method. From the results, we draw the following insights: (1) The importance weight decreases steadily throughout training, indicating that NoisyGRPO initially relies more on prior estimation driven by noise level, but gradually shifts to relying on correctness-based reward signals as training progresses. (2) NoisyGRPO consistently produces shorter completions than GRPO while achieving better performance, suggesting that its supervision mechanism over the CoT process encourages more concise and accurate reasoning steps. (3) Compared to GRPO, NoisyGRPO yields greater improvements on tasks requiring fine-grained visual understanding, such as Fine-grained Perception and Instance Reasoning, but performs slightly worse on knowledge-intensive tasks like Scene & Technology. This indicates that NoisyGRPO primarily enhances vision-language alignment in CoT reasoning, rather than boosting knowledge retrieval or external fact-based reasoning capabilities.
Figure 3: Performance over iteration and training statistics. We report the evaluation results of NoisyGRPO-3B on the MMStar benchmark to demonstrate the comprehensive capabilities of the MLLM. For both Importance Weight and Completion Length, the shaded regions represent the variance across samples.
4.6 Ablation Study
Method Design
To investigate the contribution of each component in our proposed NoisyGRPO framework, we conduct an ablation study on the comprehensive MMStar [mmstar] benchmark. As illustrated in Figure 4, we compare the performance trajectories of four variants throughout training. The baseline GRPO [GRPO] corresponds to the original policy optimization algorithm introduced in DeepSeek-R1 [DeepSeekR1]. The variant GRPO with Noise Injection incorporates our proposed noise-injected exploration policy on top of GRPO. Naive NoisyGRPO refers to a simplified version of NoisyGRPO, where the dynamic weighting mechanism between the observation and prior is removed—specifically, in Eq. 9, we set $\sigma_{s}^{2}=\sigma_{n}^{2}$, thereby eliminating the adaptive fusion. From the results, we observe that naively injecting noise to perturb the visual inputs degrades policy learning performance, indicating that a well-designed mechanism is required to mitigate the adverse effects of noise. Furthermore, the NoisyGRPO with dynamic prior-observation fusion consistently outperforms other variants, demonstrating the effectiveness of this design in enhancing the reasoning capability of MLLMs during bootstrapping.
Hyperparameter Sensitivity
To assess the sensitivity of NoisyGRPO to its hyperparameters, we conduct experiments analyzing the impact of two key parameters in the Bayesian Advantage Estimation module. As defined in Eq. 8, $\alpha$ controls the confidence assigned to the observation, while $\gamma$, defined in Eq. 9, serves as a scale hyperparameter that modulates how strongly the variability of semantic rewards influences the prior uncertainty. As shown in Figure 4, NoisyGRPO exhibits relatively low sensitivity to $\alpha$ compared to $\gamma$, indicating that the bias introduced by the embedding-based reward function is smaller than that resulting from inaccurate prior estimation due to noise injection. We further extend the analysis to the reward-temperature parameter $\tau$. As shown in Table LABEL:tab:hyperparam, NoisyGRPO maintains stable performance across a broad range of values for both hyperparameters on MMStar, AMBER, and MMERealworld benchmarks, consistently outperforming the vanilla GRPO baseline. These results confirm the robustness of NoisyGRPO and demonstrate that reasonable hyperparameter choices can be made without extensive tuning.
Figure 4: Performance of Ablation Study. We report the average performance on the MMStar benchmark across the following ablation settings: (a) Ablation of core design choices in NoisyGRPO; (b) Sensitivity to the hyperparameter $\alpha$ controlling observation confidence; (c) Sensitivity to the hyperparameter $\gamma$ modulating prior variance adaptation; (d) Comparison of different noise distributions used in the noise-injected exploration policy.
Noise Distribution in Noise Injection
To investigate the effect of different noise distributions on policy learning in NoisyGRPO, we experiment with three types of injected noise during training. Specifically, we compare: (1) Uniform noise, which is the default choice in our final method; (2) Gaussian noise sampled from a normal distribution with mean 0 and variance 0.1; and (3) Fixed noise with a constant value of 0.5. As shown in Figure 4, we observe that noise sampled from a uniform distribution consistently leads to more stable and improved MLLM performance. In contrast, both Gaussian and fixed noise result in unstable training dynamics. We attribute this to the fact that uniform noise provides a broader and more balanced exploration space, while our proposed Bayesian Advantage Estimation effectively mitigates the adverse impact of high-variance noise during policy optimization.
The Shapiro-Wilk test assesses the null hypothesis that the sample is drawn from a normal distribution. Although the p-value is very small (due to the large sample size), the $W$ value remains reasonably high, suggesting the distribution retains approximate Gaussian characteristics.
  • Kolmogorov-Smirnov Test: $D=0.190$, $p\text{-value}=3.529\times 10^{-155}$ The Kolmogorov-Smirnov test measures the maximal deviation $D$ between the empirical distribution and a reference Gaussian. Despite the low p-value (again, expected in large-scale settings), the moderate $D$ value supports the plausibility of a Gaussian assumption.
    Taken together with the results from the prior-observation residual analysis, these findings provide empirical support for the validity of our Gaussian assumption in modeling both the prior and the observation noise.
  • Appendix C Visualization
    Figure 9: Illustration of "Answer Correctness is a Partial Observation"
    Answer Correctness is a Partial Observation
    One of our key motivations stems from the fact that the correctness of the final answer cannot fully reflect the overall rollout quality, especially in complex multi-step reasoning settings like Chain-of-Thought (CoT). To qualitatively illustrate this phenomenon, we provide a concrete example in Figure 9. In this example, NoisyGRPO samples multiple rollouts under different noise levels. Interestingly, all sampled answers are correct with respect to the final answer, yet the quality of the intermediate CoT reasoning varies noticeably. This demonstrates that relying solely on answer correctness as a training or evaluation signal may overlook significant variations in reasoning quality. Hence, a more nuanced reward structure, as introduced in our method, is necessary to capture the semantics of the full reasoning trajectory.
    Figure 10: Illustration of Training rollouts and trajectory reward.
    Figure 11: Illustration of Training rollouts and trajectory reward.
    Weight in Bayesian Estimation
    In this section, we qualitatively illustrate how the training process dynamically balances the weights between the prior and observation in our Bayesian estimation framework. Specifically, the weight assigned to the observation is defined as
    $\frac{\sigma_{s}^{2}}{\sigma_{n}^{2}+\sigma_{s}^{2}},$
    where $\sigma_{n}^{2}$ and $\sigma_{s}^{2}$ denote the variances of the noise prior and semantic observation, respectively.
    We first demonstrate an extreme case where all the answers in the sampled trajectories are incorrect. In this situation, vanilla GRPO fails to provide meaningful training signals, as the semantic correctness reward becomes uniformly zero. However, as shown in Figure 10, NoisyGRPO adaptively reduces the weight of the observation (correctness score) to zero, allowing the noisy prior alone to contribute to the training signal. This enables the policy model to continue learning from the noise level’s implication on CoT quality, even in the absence of correct final answers.
    In contrast, when the correctness scores exhibit high variance—meaning some answers are correct and others are not, as illustrated in Figure 11—NoisyGRPO assigns a balanced weight between the prior and the observation. This allows the model to leverage both the semantic correctness and the injected noise level to assess the overall trajectory quality more accurately.
    These examples highlight the importance of dynamic weighting in our Bayesian estimation, enabling stable and informative policy updates across varying rollout qualities.
    Figure 12: Comparison of rollouts generation between vanilla GRPO and NoisyGRPO.
    Rollouts Sampling Comparison with Vanilla GRPO
    In this section, we compare the policy exploration characteristics of vanilla GRPO and our proposed NoisyGRPO during rollout sampling. Specifically, we examine the differences between two exploration strategies: one that solely relies on temperature sampling to generate multiple rollouts, and another that introduces noise injection to the inputs to diversify the rollouts.
    As illustrated in Figure 12, we observe that rollouts sampled by vanilla GRPO tend to exhibit high similarity to each other. This limited diversity constrains the scope of policy exploration and may lead to suboptimal learning. In contrast, NoisyGRPO, by injecting structured noise during rollout generation, encourages exploration of more diverse reasoning paths and CoT structures. This diversity allows the model to better capture the relationship between reasoning quality and task performance, ultimately leading to more robust policy improvement.
    Figure 13: CoT Inference comparison between GRPO and Noisy.
    Inference Comparison with Vanilla GRPO
    In this section, we present a qualitative visualization comparing the CoT inference behavior of vanilla GRPO and our proposed NoisyGRPO. As shown in Figure 13, we observe that rollouts generated by NoisyGRPO tend to have shorter chains of thought while maintaining comparable answer accuracy. This suggests that NoisyGRPO achieves more efficient CoT reasoning by avoiding unnecessary elaboration.
    Moreover, we find that in some cases, NoisyGRPO generates an intermediate question within the CoT, which serves as a refinement or decomposition of the original input question. This strategy allows the model to clarify its internal reasoning path and arrive at a more accurate final answer. In contrast, vanilla GRPO often produces longer, redundant reasoning steps that do not necessarily contribute to better accuracy.
    These observations indicate that incorporating noisy priors during training encourages the policy model to develop more concise and targeted reasoning strategies during inference.
    Comparison with Rule-Based Reward
    To validate our choice of using the text embedding model SBERT as the reward model, we compare it with a popular rule-based reward approach. Specifically, we adopt the average F1 score of ROUGE‑1, ROUGE‑2, and ROUGE‑L as the reward. The results are summarized in Table 5. As shown, SBERT consistently outperforms the rule-based reward in our experimental setting, likely due to its ability to capture semantic similarity beyond exact token overlap, which is especially important for evaluating open-ended answers where lexical variance is common.
    Table 5: Comparison of SBERT-based reward vs. rule-based ROUGE reward on MMStar, AMBER, and MMERealworld metrics.
    Method MMStar AMBER MMERW
    CP. FP. IR. LR. ST. MA. Avg. Cs ↓ Cov. ↑ Hal. ↓ Cog. ↓ F1 ↑ Acc
    Rouge reward 64.9 49.5 64.8 51.3 38.0 52.6 53.5 8.0 66.4 46.5 3.4 89.3 43.6
    SBERT reward 69.5 53.2 66.6 60.7 37.1 62.3 58.2 6.6 67.7 44.3 3.4 90.3 44.0
    Appendix D Scalability Analysis
    To evaluate the scalability of NoisyGRPO with respect to training data size, we conduct experiments using subsampled datasets containing 3k and 6k examples, compared to the original 13k dataset. Specifically, we randomly sample 3k and 6k subsets from the full training corpus and train both GRPO and NoisyGRPO under identical training steps to ensure a fair comparison.
    As shown in Table 6, NoisyGRPO maintains consistent gains over the GRPO baseline across all dataset sizes, demonstrating its robustness and scalability when exposed to larger data volumes. Furthermore, performance improvements are observed as the amount of training data increases, confirming that NoisyGRPO effectively leverages additional data to enhance policy learning. These results highlight that NoisyGRPO is a scalable and data-efficient optimization framework, capable of leveraging additional training samples to further enhance multimodal reasoning performance without loss of stability.
    Table 6: Scalability analysis of NoisyGRPO. Performance comparison between GRPO and NoisyGRPO trained with subsampled datasets (3k and 6k samples). NoisyGRPO consistently outperforms the baseline under all data scales, and its performance improves with larger datasets. Best results in each column are highlighted.
    Method / Size MMStar AMBER MMERW
    CP. FP. IR. LR. ST. MA. Avg. Cs ↓ Cov. ↑ Hal. ↓ Cog. ↓ F1 ↑ Acc
    GRPO 3k 66.4 44.6 59.7 53.8 35.1 56.7 52.7 7.1 66.1 46.9 3.8 88.9 38.9
    NoisyGRPO 3k 66.7 51.8 61.2 56.4 35.3 62.3 55.6 6.7 63.5 36.8 2.2 89.9 44.5
    GRPO 6k 67.2 45.6 58.7 53.2 40.2 54.1 53.2 7.0 67.3 48.8 4.7 88.3 38.1
    NoisyGRPO 6k 69.1 52.2 64.3 54.7 39.7 59.5 56.6 6.4 65.4 40.2 3.1 90.1 44.1